### v1

@@ Abstract
We present an implementation and a brief experimental analysis of the deterministic algorithm proposed by Duan et al. (2025) for the Single-Source Shortest Path (SSSP) problem, which achieves the best known asymptotic upper bound in the comparison-addition model, with running time $O(m\log^{2/3}n)$. We provide a faithful C++ implementation of this algorithm, following all structural details described in the original paper, and compare its empirical performance with the classical Dijkstra’s algorithm using binary heaps. The experiments were conducted on both synthetic sparse random graphs and real-world road network instances from the DIMACS benchmark. Our results show that, despite its superior asymptotic complexity, the new algorithm presents significantly larger constant factors, making Dijkstra’s algorithm faster for all tested sparse graph sizes, including instances with tens of millions of vertices. Our implementation achieves $O(m\log^{2/3}n)$ expected time, due to the use of hash tables, and some possibilities for making it worst-case are being considered. (This is a ongoing work.)
@@
Single Source Shortest Path, Dijkstra’s Algorithm, Graph, Algorithms, Experimental Analysis, Time Complexity
@@ 1 Introduction
This work aims to empirically evaluate the deterministic algorithm proposed by duan2025breaking and compare its practical performance with Dijkstra’s classical algorithm. We implement both algorithms in C++ (Dijkstra’s algorithm using a binary heap, and Duan’s algorithm faithfully following the implementation details described in their paper).
@@
To the best of our knowledge, this is the first work to present a faithful implementation of Duan’s algorithm and provide a practical performance analysis on random and road networks up to approximately 10M vertices.
@@
The algorithm uses two parameters, $k$ and $t$, and is based on a divide-and-conquer approach over the vertex set. The vertex set is recursively divided into $2^{t}$ roughly equal parts across $O((\log n)/t)$ recursion levels, until reaching the base case where a subproblem contains a single vertex $x$. At the base case, Dijkstra’s algorithm is executed to compute the shortest distances from $x$ to its $k$ nearest vertices.
@@
3 Constant Analysis
@@ 4 Implementation Details
We provide implementations for both algorithms discussed in this work. All implementations are written in C++20, compiled with g++ using the -O3 optimization flag. The code is publicly available at github.com/lcs147/bmssp.
Dijkstra’s algorithm was implemented using the binary heap provided by the C++ standard library (priority_queue).
The algorithm of duan2025breaking was implemented faithfully according to the description in their paper. The original algorithm is designed to run in $O(m\log^{2/3}n)$ worst-case time. Our implementation achieves $O(m\log^{2/3}n)$ expected time, due to the use of hash tables for certain set operations (such as union, membership tests, and duplicate removal). The paper does not specify the data structures for these operations, and we found no alternative that would preserve both the $O(m)$ memory usage and the $O(m\log^{2/3}n)$ time complexity. Therefore, hash tables were chosen as a practical solution for implementing these operations efficiently.
@@ 5 Experiments
  • Sparse Random Graphs: randomly generated graphs with sizes $2^{7},2^{8},\dots,2^{24}$. These graphs have a mean out-degree of 3 and a maximum out-degree of 4. The vertices were numbered from 1 to $n$, and it is guaranteed that vertex 1 can reach all other vertices. The code used to generate these instances is available in the repository.
@@ 6.1 Sparse Random Graphs
Figure 1 shows the average execution time (on logarithmic scale) for the randomly generated graphs. Graphs with up to $2^{24}$ vertices were tested, which was the maximum size the BMSSP algorithm could handle due to the machine memory limitations.
Figure 1: Execution time (average over multiple runs) for randomly generated sparse graphs of increasing size, plotted on a log-log scale.
Figure 2 illustrates the ratio of execution times between the BMSSP and Dijkstra algorithms on random graphs. For the smallest graph size (128 vertices), BMSSP is substantially slower, measuring approximately 22 times slower than Dijkstra. As the graph size increases, however, this performance disparity significantly narrows, with the ratio dropping to approximately seven. This decreasing ratio indicates that BMSSP becomes relatively more efficient as the input size grows. Overall, across all random graph instances, BMSSP maintains an average execution time approximately nine times slower than Dijkstra.
@@ 6.2 Road Graphs
Table 1 reports execution times for the 12 road network instances, and Figure 4 illustrates the growth of execution time with increasing graph size. Dijkstra’s algorithm was on average approximately seven times faster than the BMSSP algorithm.
While this ratio (seven) is lower than the ratio observed in the random graph experiments (nine), this result is expected. The road network instances start significantly larger, with the smallest network already exceeding 200,000 vertices (compared to 128 vertices in the smallest random graph). This larger scale provides an environment where the overhead of the BMSSP algorithm is relatively less impactful, thus reducing the performance gap between it and Dijkstra’s algorithm.
@@
Instance $n$ $m$ Time (ms) Ratio
@@
New York City 264,346 733,846 32.490 250.030 7.696
San Francisco Bay Area 321,270 800,172 38.803 297.297 7.662
Colorado 435,666 1,057,066 52.174 416.865 7.990
Florida 1,070,376 2,712,798 127.588 958.971 7.516
Northwest USA 1,207,945 2,840,208 155.193 1067.577 6.879
Northeast USA 1,524,453 3,897,636 228.756 1450.339 6.340
California and Nevada 1,890,815 4,657,742 257.454 1719.994 6.681
Great Lakes 2,758,119 6,885,658 376.873 2674.383 7.096
Eastern USA 3,598,623 8,778,114 553.878 3718.830 6.714
Western USA 6,262,104 15,248,146 1013.046 6903.403 6.815
Central USA 14,081,816 34,292,496 3346.156 18195.013 5.438
Full USA 23,947,347 58,333,344 4562.473 29345.319 6.432
@@ 6.3 General analysis
If we assume that the constant of our BMSSP implementation is seven times higher than Dijkstra’s ($c_{2}/c_{1}=7$), a ratio derived from the road network experiment, then the same analysis used in Section 3 predicts that our implementation will only be faster when the graph size surpasses $10^{103}$.
@@ 7 Concluding Remarks
Our experiments show that Dijkstra’s algorithm remains the preferred choice for graphs with up to ten million vertices, covering most practical applications. Also, we predict that Dijkstra’s should be the preferred choice for graphs much bigger than that. (As discussed in Section 3, its large hidden constants limit its practical efficiency, despite its superior asymptotic bound.)

### v2

@@ Abstract
We present an implementation and a brief experimental analysis of the deterministic algorithm proposed by Duan et al. (2025) for the Single-Source Shortest Path (SSSP) problem, which achieves the best known asymptotic upper bound in the comparison-addition model, with running time $O(m\log^{2/3}n)$. We provide a faithful C++ implementation of this algorithm, following all structural details described in the original paper, and compare its empirical performance with the classical Dijkstra’s algorithm using binary heaps. The experiments were conducted on both synthetic sparse random graphs and real-world road network instances from the DIMACS benchmark. Our implementation achieves the proposed running time in the worst case. However, our results show that, despite the new algorithm’s superior asymptotic complexity, its large constant factors significantly affect its practical performance, making Dijkstra’s algorithm faster for all tested sparse graph sizes, including instances with tens of millions of vertices. (This is a work in progress.)
@@
Algorithms, Dijkstra’s Algorithm, Experimental Analysis, Heap and Hash Data Structures, Single Source Shortest Path, Time Complexity.
@@ 1 Introduction
This work aims to empirically evaluate the deterministic algorithm proposed by duan2025breaking and compare its practical performance with Dijkstra’s classical algorithm. We implement both algorithms in C++ (Dijkstra’s algorithm using a binary heap, and duan2025breaking algorithm faithfully following the implementation details described in their paper).
@@
During the development of this work, we also became aware that makowski2025bmsspy are developing a Python package that aims to faithfully implement the algorithm proposed by duan2025breaking. They did experiments on grid and geospatial networks to compare their implementations with other Python packages which implement Dijkstra’s algorithm.
To the best of our knowledge, this is the first work to present a faithful implementation of the duan2025breaking algorithm and to provide a practical performance analysis on random and road networks up to approximately 10 million vertices.
@@
The algorithm uses two parameters, $k$ and $t$, and is based on a divide-and-conquer approach over the vertex set. The vertex set is recursively divided into $2^{t}$ roughly equal parts across $O((\log n)/t)$ recursion levels, until reaching the base case where a subproblem contains a single vertex $x$. In the base case, Dijkstra’s algorithm is executed to compute the shortest distances from $x$ to its $k$ nearest vertices.
@@
3 Asymptotic versus Empirical Analysis under Big Constants
@@ 4 Implementation Details
We provide implementations for both algorithms discussed in this work. All implementations are written in C++20, compiled with g++ using the -O3 optimization flag. The code is publicly available at https://github.com/lcs147/bmssp.
The algorithm proposed by duan2025breaking requires that all simple paths have distinct lengths. To satisfy this requirement in a general setting, we must break ties consistently when paths have equal length. As described in the original paper, we represent the distance to a vertex $v$ as the tuple
$\hat{d}[v]=\big(\text{length}[v],\ \text{hops}[v],\ v,\ \text{pred}[v]\big).$
Here, $\text{length}[v]$ is the total weight of the current shortest path to $v$, $\text{hops}[v]$ is the number of edges in this path, and $\text{pred}[v]$ is the predecessor of $v$ in this path. Paths are compared using the lexicographic order on these tuples, which guarantees a unique ordering even if their length values are identical.
Dijkstra’s algorithm was implemented using the binary heap provided by the C++ standard library (std::priority_queue). We selected this data structure because binary heaps are often more efficient in practice than Fibonacci heaps (cherkassky1996shortest).
Our implementation closely follows the algorithm proposed by duan2025breaking. We implemented all three subroutines and the core data structure as specified, ensuring that the theoretical worst-case time complexity of $O(m\log^{2/3}n)$ is preserved. Below, we list some implementation details:
Additional memory usage
The block-based data structure in the original paper stores at most $N$ key-value pairs. It must support finding any key and its associated data in $O(1)$ time.
Instead of using a hash table to find the keys (which provides expected $O(1)$ operations), we employ global direct-address tables (DATs) of size $O(n)$, where $n$ is the number of vertices. Naively, each recursive call of BMSSP (Algorithm 3) would require its own DAT, leading to an infeasible number of tables. However, due to the sequential nature of the divide-and-conquer recursion, we can reuse DATs across calls. Specifically, we maintain only one DAT per recursion level, as recursive calls at the same level are processed sequentially and can share the same memory. With $O(\log^{1/3}n)$ recursion levels and a DAT of size $\Theta(n)$ per level, the total space complexity becomes $\Theta(n\log^{1/3}n)$.
Set operations
To guarantee $O(1)$ worst-case operations for set operations (such as union, membership tests, and duplicate removal), we consistently use global direct-address tables (DATs) throughout our implementation. However, unlike in the bucket data structure, the other DATs do not increase the asymptotic space complexity.
FindPivots forest building.
In the FindPivots subroutine, we do not explicitly build the directed forest $F$. Since we only need to identify roots whose trees contain at least $k$ vertices, we maintain an array $\text{root}[v]$ that stores the root of the tree to which vertex $v$ belongs. Whenever an edge $(u,v)$ is relaxed, we update $\text{root}[v]:=\text{root}[u]$. This approach allows us to efficiently determine the final set of pivots without the overhead of building the full forest and traversing any vertex.
Disjointness of $U_{i}$ sets
Experimentally, we encountered a subtle discrepancy with Remark 3.8 in duan2025breaking, which asserts that the sets $U_{i}$ in a recursive call of BMSSP (Algorithm 3) are disjoint. This property may fail. Consider a vertex $x$ that enters the data structure $D$ with an initial distance $d_{\text{old}}$. If $x$ is later marked as complete by a recursive call with a shorter distance $d_{\text{new}}$ (so that $x\in U_{i}$) but is never extracted from $D$, it remains in $D$ with the outdated key $d_{\text{old}}$. When $x$ is eventually extracted (still with key $d_{\text{old}}$), it becomes a frontier vertex, causing its shortest‑path subtree to be rediscovered with the correct distance $d_{\text{new}}$. Consequently, $x$ and its subtree may be assigned to another set $U_{j}$, violating disjointness.
We restore the disjointness property by removing a vertex from $D$ as soon as it is marked complete. Specifically, before line 15 of subroutine BMSSP (Algorithm 3), we perform $D.\mathtt{erase}(u)$. Assuming $\mathtt{erase}$ has complexity $O(\max\{1,\log(N/M)\})$ when the element is present and $O(1)$ otherwise, the asymptotic complexity of the algorithm is preserved. With this modification, the implementation satisfies the disjointness claim of Remark 3.8.
@@ 5 Experiments
  • Sparse Random Graphs: randomly generated graphs with sizes $2^{7},2^{8},\dots,2^{25}$. These graphs have a mean out-degree of 3 and a maximum out-degree of 4. The vertices were numbered from 1 to $n$, and it is guaranteed that vertex 1 can reach all other vertices. The code used to generate these instances is available in the repository.
@@ 6.1 Sparse Random Graphs
Table 1 reports execution times for sparse random graphs with mean out-degree of 3. Experiments were conducted on graphs with up to $2^{25}$ vertices, the maximum size that could be accommodated within the available memory. Attempts to process a graph with $2^{26}$ vertices resulted in memory exhaustion. The scaling of average execution time (on a logarithmic scale) across graph sizes is illustrated in Figure 1.
Figure 1: Execution time for random sparse graphs of increasing size, on a log-log scale.
Figure 2 illustrates the ratio of execution times between the BMSSP and Dijkstra algorithms on random graphs. For the smallest graph size ($128$ vertices), BMSSP is substantially slower, measuring approximately $24$ times slower than Dijkstra. However, as the graph size increases, this performance disparity significantly narrows, with the ratio dropping to approximately $3$ for some graphs. This decreasing ratio indicates that BMSSP becomes relatively more efficient as the input size grows. Overall, across all random graph instances, BMSSP maintains an average execution time approximately four times higher than Dijkstra.
Table 1: Execution times (in milliseconds) of our implementations of Dijkstra and Duan et al. algorithms on random sparse graphs.
Instance $n$ $m$ Time (ms) Ratio
dljkstra1959note duan2025breaking
random128D3 $2^{7}$ $3\times 2^{7}$ $0.006\pm 0.003$ $0.148\pm 0.012$ 24.667
random256D3 $2^{8}$ $3\times 2^{8}$ $0.013\pm 0.004$ $0.163\pm 0.010$ 12.538
random512D3 $2^{9}$ $3\times 2^{9}$ $036\pm 0.005 0.0$$087\pm 0.011$ 7.194
random1024D3 $2^{10}$ $3\times 2^{10}$ $0.087\pm 0.006$ $0.575\pm 0.021$ 6.609
random2048D3 $2^{11}$ $3\times 2^{11}$ $0.192\pm 0.006$ $1.175\pm 0.028$ 6.120
random4096D3 $2^{12}$ $3\times 2^{12}$ $0.431\pm 0.017$ $2.244\pm 0.038$ 5.206
random8192D3 $2^{13}$ $3\times 2^{13}$ $1.036\pm 0.081$ $634\pm 0.058 4.4$73
random16384D3 $2^{14}$ $3\times 2^{14}$ $2.199\pm 0.040$ $8.639\pm 0.410$ 3.929
random32768D3 $2^{15}$ $3\times 2^{15}$ $4.769\pm 0.068$ $20.455\pm 0.539$ 4.289
random65536D3 $2^{16}$ $3\times 2^{16}$ $11.237\pm 0.143$ $44.527\pm 0.413$ 3.963
random131072D3 $2^{17}$ $3\times 2^{17}$ $28.035\pm 0.654$ $95.711\pm 1.126$ 3.414
random262144D3 $2^{18}$ $3\times 2^{18}$ $73.609\pm 0.903$ $237.379\pm 2.058$ 3.225
random524288D3 $2^{19}$ $3\times 2^{19}$ $174.674\pm 2.412$ $631.082\pm 1.475$ 3.613
random1048576D3 $2^{20}$ $3\times 2^{20}$ $418.830\pm 2.058$ $1450.863\pm 0.565$ 3.464
random2097152D3 $2^{21}$ $3\times 2^{21}$ $946.968\pm 7.659$ $3226.209\pm 8.977$ 3.407
random4194304D3 $2^{22}$ $3\times 2^{22}$ $2012.560\pm 3.322$ $8218.205\pm 13.655$ 4.083
random8388608D3 $2^{23}$ $3\times 2^{23}$ $4494.410\pm 5.503$ $15646.055\pm 11.004$ 3.481
random16777216D3 $2^{24}$ $3\times 2^{24}$ $9998.167\pm 6.375$ $34104.239\pm 188.505$ 3.411
random33554432D3 $2^{25}$ $3\times 2^{25}$ $21765.062\pm 14.522$ $85382.069\pm 49.101$ 3.923
@@ 6.2 Road Graphs
Table 2 reports execution times for the 12 road network instances, and Figure 4 illustrates the growth of execution time with increasing graph size. The results show that Dijkstra’s algorithm consistently outperforms the BMSSP algorithm by an average factor of four across all twelve road network instances. The performance ratio remains remarkably stable, ranging only from $3.67$ to $4.89$, despite graph sizes varying over two orders of magnitude.
The observed four-fold performance ratio is similar to the ratio previously measured on random graphs. Dijkstra maintains a clear advantage; the performance gap seems stable as the graph size increases.
@@
Instance $n$ (approx.) $m$ (approx.) Time (ms) Ratio
@@
New York City $2.6\times 10^{5}$ $7.3\times 10^{5}$ $33.694\pm 0.266$ $138.444\pm 0.731$ 4.109
San Francisco Bay Area $3.2\times 10^{5}$ $8.0\times 10^{5}$ $39.672\pm 0.535$ $163.748\pm 1.228$ 4.128
Colorado $4.3\times 10^{5}$ $1.0\times 10^{6}$ $51.731\pm 0.356$ $253.033\pm 1.556$ 4.891
Florida $1.0\times 10^{6}$ $2.7\times 10^{6}$ $126.690\pm 1.733$ $546.35\pm 5.189$ 4.313
Northwest USA $1.2\times 10^{6}$ $2.8\times 10^{6}$ $153.513\pm 1.658$ $621.822\pm 2.873$ 4.051
Northeast USA $1.5\times 10^{6}$ $3.8\times 10^{6}$ $228.126\pm 1.654$ $836.506\pm 3.755$ 3.667
California and Nevada $1.8\times 10^{6}$ $4.6\times 10^{6}$ $258.069\pm 2.587$ $1006.176\pm 2.887$ 3.899
Great Lakes $2.7\times 10^{6}$ $6.8\times 10^{6}$ $378.239\pm 3.004$ $1626.529\pm 3.147$ 4.300
Eastern USA $3.5\times 10^{6}$ $8.7\times 10^{6}$ $541.000\pm 3.923$ $2413.533\pm 4.204$ 4.461
Western USA $6.2\times 10^{6}$ $1.5\times 10^{7}$ $1012.382\pm 2.023$ $549.724\pm 9.840 4.4$94
Central USA $1.4\times 10^{7}$ $4\times 10^{7} 33$$355.727\pm 12.259$ $14104.518\pm 21.586$ 4.203
Full USA $2.3\times 10^{7}$ $5.8\times 10^{7}$ $4566.797\pm 4.672$ $21261.355\pm 10.304$ 4.656
@@ 6.3 General analysis
If we assume that the constant of our BMSSP implementation is four times higher than Dijkstra’s ($c_{2}/c_{1}=4$), a ratio derived from the road network experiment, then the same analysis used in Section 3 predicts that our implementation will only be faster when the graph size surpasses $10^{19}$.
@@ 7 Concluding Remarks
Our experiments show that Dijkstra’s algorithm remains the preferred choice for graphs with up to ten million vertices, covering most practical applications. Furthermore, we predict that Dijkstra’s algorithm should be the preferred choice for graphs significantly larger than that.
@@
Several avenues for further investigation remain open. First, we plan to refine the BMSSP implementation by replacing its current linear-time selection routine (blum_time_1973) with a more efficient deterministic selection algorithm that offers better constant factors (alexandrescu2017fast). This modification may reduce the algorithm’s overhead and narrow the performance gap observed in our experiments.
Additionally, we will expand our experimental analysis in several directions. We intend to evaluate both algorithms on structured grid graphs, on which the Bellman-Ford algorithm has previously been shown to outperform Dijkstra (cherkassky1996shortest). We also plan to conduct systematic comparisons between expected-case and worst-case performance, implement a Fibonacci heap version of Dijkstra’s algorithm as an additional baseline, and test the impact of the degree-normalization preprocessing step required by the duan2025breaking algorithm to ensure constant out-degree.
Note
This is a work in progress.